Fermi National Laboratory


Accelerator Update

The integrated luminosity for last week was 11.30 pb-1.

Monday February 16
The day shift began with Operations monitoring store 3237, with the Antiproton Source (Pbar) stacking, and with a Main Injector (MI) expert conducting slow-acceleration studies.

At 10:16 AM, Operation reported that experts were conducting studies in Recycler (RR), and MI and Switchyard (SY). They also reported that MiniBooNE was taking beam.

Pbar experts halted stacking and began a stacking study at 11:30 AM followed by a reverse proton study period.

At 1:03 PM, Linac experts brought the I- Source vacuum up to normal atmosphere to conduct maintenance.

A MI expert began a slip-stacking study at 2:30 PM.

A problem with Linac corrector magnets shut off beam at 2:56 PM for about ten minutes.

A Booster power supply (EAPS) tripped off at 3:47 PM. Experts found a tripped 480V breaker that took down quite a few systems. At the same time a Linac RF station (KRF1) tripped off. A Linac expert had KRF1 back on by 4 PM.

Booster technicians completed their repairs and beam returned to MiniBooNE at 6:08 PM. Pbar studies resumed at 6:17 PM.

SY, Pbar, and RR experts completed their studies by 8:40 PM,

Pbar resumed stacking at 9 PM. A periodic problem with antiproton production and stack rate required operators to call in experts for help.

Operations terminated store 3237 at 10 PM. Tevatron (TeV) experts began aperture scans at 10:24 PM. The TeV suffered an abort at 11:18 PM during the studies. Everything reset okay and the studies continued.

Tuesday February 17
The midnight shift began with TeV experts conducting studies, with Pbar experts investigating stacking and production problems, and with MiniBooNE taking beam.

At 12:22 AM, the TeV suffered another abort. Everything reset okay and the studies continued.

At 1:23 AM, the TeV studies ended and Operations put the TeV into shot setup.

Operations established store 3240 at 4:47 AM with an initial luminosity of 56.91E30. Pbar resumed stacking.

Pbar stopped stacking at 7 AM and setup for shots to RR, which began at 7:37 AM.

Pbar resumed stacking at 8:47 AM.

At 12:50 PM, a MI expert began more slipstacking studies.

D-zero (D0) experimenters reported at 1:22 PM that their main magnet quenched. They recovered the magnet and resumed taking data at 2:21 PM.

Wednesday February 18
The midnight shift began with Operations monitoring store 3240, with Pbar stacking, and with MiniBooNE taking beam.

At 4:33 AM, Operations reported that RR had a cooling problem and lost a lot of their stash.

MI suffered a vacuum burst at 7:04 AM. Everything reset okay.

Today's Plans
The plans for today are to allow TeV experts to conduct end-of-store studies with the remains of store 3240, followed by a short access into the TeV later this morning.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/18/2004   email Fermilab